Suppose you have predicted the following returns for Apple Inc. stock in three possible states of...

  1. Suppose you have predicted the following returns for Apple Inc. stock in three possible states of nature. What is the expected return to Apple Inc.?

State               Probability of State                            Apple Inc. Returns

Boom                    0.4                                                                             - 0.12

Normal                0.3                                                                             0.06

Bust                                                                                                             0.27
Answer #1

Expected Return is weighted average return of the stock

Expected Return E(R) = P1 * R1 + P2 * R2 + P3 * R3

P3 = 1 - P1 - P2 = 1 - 0.4 - 0.3 = 0.3

Expected return E (R) = 0.4 * -12% + 0.3 * 6% + 0.3 * 27% = 5.1%